2. The Andersonian Consulting Company currently has five employees. The best-case
situation is for the employees to be at 70 percent utilization. At 160 hours per month and a
billing rate of $55 per hour, what is the maximum revenue that can be achieved?

3. Assume that a lawn service company collected $127,000 last month and that it charges a
typical customer $100 per month for its service. This month the revenue is $134,000 and
the sales department claims to have added 300 new customers.
a. How many customers did the company actually have and collect payments from last month?
b. How many paying customers does the company have this month?
c. How many customers probably disconnected from the company’s service in the past month?
d. What is the disconnect rate?
Answer #1

2: Maximum revenue = No of employees*Utilization rate * Billing rate per hour * No of hours

= 5*70%*55*160

= 30800

3: a: Customers that the company actually have and collect payments from last month = 127000/100 = 1270

b: Paying customers this month= 134000/100 = 1340

c: No of disconnected customers = 1270 last month +300 new added -1340 actual paying

= 230

d:Disconnect rate = 230/1270 = 18.11%